### v1

# Mechanistic Understanding and Mitigation of Language Confusion in English-Centric Large Language Models

Abstract
Language confusion—where large language models (LLMs) generate unintended languages against the user’s need—remains a critical challenge, especially for English-centric models. We present the first mechanistic interpretability (MI) study of language confusion, combining behavioral benchmarking with neuron-level analysis. Using the Language Confusion Benchmark (LCB), we show that confusion points (CPs)—specific positions where language switches occur—are central to this phenomenon. Through layer-wise analysis with TunedLens and targeted neuron attribution, we reveal that transition failures in the final layers drive confusion. We further demonstrate that editing a small set of critical neurons, identified via comparative analysis with multilingual-tuned models, substantially mitigates confusion without harming general competence or fluency. Our approach matches multilingual alignment in confusion reduction for most languages and yields cleaner, higher-quality outputs. These findings provide new insights into the internal dynamics of LLMs and highlight neuron-level interventions as a promising direction for robust, interpretable multilingual language modeling. Code and data will be released upon publication.
Mechanistic Understanding and Mitigation of Language Confusion
in English-Centric Large Language Models
Ercong Nie1,2   Helmut Schmid1   Hinrich Schütze1,2 1Center for Information and Language Processing (CIS), LMU Munich, Germany 2 Munich Center for Machine Learning (MCML), Germany nie@cis.lmu.de
1 Introduction
Figure 1: Language Confusion in LLMs. (a) An example of the language confusion phenomenon. (b) Visualization of internal model dynamics using TunedLens, highlighting how the confusion point emerges during generation. (c) Benchmarking results of three Llama models on the LCB benchmark across 5 languages.
Current Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023), PaLM 2 (Anil et al., 2023), and Llama 3 (Grattafiori et al., 2024), have demonstrated exceptional linguistic competence across a wide range of complex tasks that require abstract knowledge and reasoning (Dong et al., 2024; Wei et al., 2022). Early LLMs were predominantly trained on massive amounts of English text data, with some limited exposure to other languages, resulting in initially constrained multilingual capabilities (Touvron et al., 2023). Recent advances, such as multilingual continued pretraining and instruction tuning, have substantially extended these models’ ability to support multiple languages (Zhu et al., 2023; Shaham et al., 2024; Kew et al., 2024; Wang et al., 2025b). As a result, contemporary English-centric LLMs have become foundational tools for multilingual communication, multilingual content generation, and cross-lingual applications (Bang et al., 2023; Ahuja et al., 2023; Asai et al., 2024). However, despite their impressive capabilities, a persistent and underexplored limitation remains: LLMs can fail to generate text in the user’s intended language, even when explicitly instructed—a phenomenon termed language confusion (Marchisio et al., 2024). Language confusion manifests as full-response, line-level, or word-level switches into unintended languages, severely undermining user experience and model reliability, especially for non-English speakers (Figure 1a).
Recent work by  Marchisio et al. (2024) provides the first systematic characterization of language confusion, introducing the Language Confusion Benchmark (LCB) and associated metrics to quantify this phenomenon across a diverse set of languages and models. Their evaluation revealed that even state-of-the-art LLMs are susceptible to language confusion, with English-centric LLMs such as Llama2, Llama3, and Mistral exhibiting particularly high rates of unintended language switching, especially in the absence of targeted multilingual alignment (Figure 1c). While  Marchisio et al. (2024) propose several mitigation strategies, including decoding adjustments, prompting techniques, and multilingual fine-tuning, these approaches remain largely surface-level, offering limited insight into the internal mechanisms that give rise to language confusion.
A key observation from prior work is the identification of confusion points—specific positions in the generation process where the model abruptly switches to an unintended language. However, the model’s internal dynamics leading to these confusion points and their causal role in language confusion remain largely unexplored. This gap is particularly salient given the parallels to human bilingual code-switching, where switch points between languages are cognitively significant as extensively studied in psycholinguistics (Solorio and Liu, 2008; Bullock and Toribio, 2009). Further discussions are provided in Appendix A.
In this work, we move beyond behavioral evaluation to open the black box of LLMs, leveraging mechanistic interpretability (MI) methods (Conmy et al., 2023; Rai et al., 2024; Saphra and Wiegreffe, 2024; Sharkey et al., 2025) to investigate the internal representations and neuron-level processes underlying language confusion. We first empirically demonstrate that confusion points are critical drivers of language confusion: targeted interventions at these points can substantially reduce confusion across languages. Building on this, we employ MI tools such as TunedLens (Belrose et al., 2023) to trace the evolution of language representations through the model’s layers, revealing that confusion typically arises from transition failures in the final layers, where latent conceptual representations are mapped to surface forms in the target language (Figure 1b). To further elucidate the mechanism, we conduct a neuron-level analysis, identifying specific neurons in the last layers whose activity is predictive of successful or failed language transitions at confusion points. Inspired by recent advances in neuron attribution and editing, we show that targeted manipulation of only 100 neurons can mitigate language confusion, offering a novel, model-internal approach to improving multilingual reliability. Our findings provide the first mechanistic account of language confusion in LLMs, bridging the gap between behavioral benchmarks and internal model dynamics. By highlighting the central role of confusion points and their neural substrates, we lay the groundwork for more robust, interpretable, and cognitively informed multilingual language models.
Our work makes the following contributions: (1) We provide the first mechanistic interpretability study of language confusion in English-centric LLMs, revealing the central role of confusion points in unintended language switching; (2) We employ layer-wise and neuron-level analyses to trace the internal dynamics leading to language confusion and identify critical late-layer neurons responsible for transition failures; (3) We propose and validate a principled neuron selection and editing strategy that effectively mitigates language confusion and preserves the model’s general competence and output quality.
Dataset Data Source Language Prompt Example
Aya Human-generated ar, en, pt, tr, zh {CJK*}UTF8gbsn 请简单介绍诗人李白的背景。
(Singh et al., 2024) Briefly introduce the poet Li Bai.
Dolly MT post-edited ar, es, fr, hi, ru Qu’est-ce qui est plus important, l’inné ou l’acquis?
(Singh et al., 2024) What is more important, nature or nurture?
Native Human-generated es, fr, ja, ko {CJK}UTF8mj 콘크리트는 뭘로 만든거야?
(Marchisio et al., 2024) What is concrete made of?
Okapi Synthetic + MT ar, en, pt, zh,it, Schreib einen Aufsatz von 500 Wörtern zum Thema KI.
(Lai et al., 2023) fr, de, id, es, vi Write a 500-word essay on AI.
Table 1: Overview and Prompt Example of the LCB Benchmark (monolingual part). The number of examples per language is 100 in each dataset.
2 Related Work
Mechanistic Interpretability Methods
Mechanistic interpretability (MI) seeks to reverse-engineer neural networks by decomposing their computations into human-understandable components (Stolfo et al., 2023; Wang et al., 2024; Men et al., 2024). A central technique in MI is the projection of intermediate representations into the vocabulary space, as implemented by tools such as LogitLens (Nostalgebraist, 2020) and TunedLens (Belrose et al., 2023), which enable researchers to track how information and predictions evolve across layers (Dar et al., 2023; Pal et al., 2023). In addition to layer-wise analysis, recent work has focused on identifying, attributing, and intervening on important neurons—those whose activations are strongly correlated with specific linguistic functions or behaviors (Bau et al., 2020; Geva et al., 2022; Yu and Ananiadou, 2024b). Methods for neuron selection and editing, as well as circuit-level analysis (Elhage et al., 2021; Wang et al., 2023), have proven effective for uncovering the internal structure underlying phenomena such as factual recall (Meng et al., 2022; Geva et al., 2023), reasoning processing (Yu and Ananiadou, 2024a), and now, as in our work, language confusion. By leveraging these MI techniques, we aim to provide a granular, causal understanding of how and why language confusion arises in multilingual LLMs, and to identify actionable intervention points for mitigation.
Model Metric ar en pt tr zh es fr hi ru ja ko de id it vi avg
Llama3 LPR 33.0 99.5 71.0 33.0 19.3 73.0 59.3 8.0 28.0 14.0 23.0 19.0 22.0 34.0 11.0 36.5
(original) Acc 33.7 99.8 74.5 37.5 23.4 77.1 64.1 15.1 28.2 17.1 23.6 23.0 27.3 39.8 14.8 39.9
Llama3 LPR 71.0 99.0 93.0 50.0 57.3 94.3 84.0 37.0 78.6 50.0 45.0 60.0 67.0 86.0 62.0 68.9
(replace) Acc 74.8 99.6 95.4 55.5 64.1 95.3 86.5 47.6 83.1 55.3 48.6 62.3 77.7 87.5 66.1 73.3
Llama3 LPR 98.3 98.5 99.0 95.8 88.8 98.3 95.9 97.0 100.0 93.5 100.0 100.0 88.8 100.0 97.9 96.8
(multilingual) Acc 98.7 99.5 99.8 96.9 93.8 99.3 96.9 97.5 100.0 95.8 100.0 100.0 94.2 100.0 97.9 98.0
Table 2: Impact of Confusion Point Replacement on Language Confusion Metrics. Line-level pass rate (LPR) and line-level accuracy for original Llama3-8B, multilingual Llama3-8B, and Llama3-8B with confusion point replacement, reported by language.
Multilingual Interpretability
Recent research has begun to probe the internal representations of English-centric and multilingual LLMs to understand how they process and transfer information across languages (He et al., 2024; Zhao et al., 2024)Wendler et al. (2024) show that models like Llama2 often rely on English as an internal pivot language and can disentangle language and conceptual representations in controlled tasks. Fierro et al. (2025) examine how mechanisms identified in monolingual contexts generalize to multilingual settings. Wang et al. (2025a) investigate the internal causes of crosslingual factual inconsistencies, revealing how MLMs transition from language-independent to language-specific processing. However, prior work has not systematically connected these internal mechanisms to language generation errors such as language confusion.
3 Revisiting Language Confusion: Benchmark Insights
3.1 Recap of Language Confusion Benchmark
The Language Confusion Benchmark (LCB) (Marchisio et al., 2024) provides a systematic framework for evaluating the ability of LLMs to generate text in the user’s intended language. The benchmark covers 15 typologically diverse languages and uses a diverse set of prompts sourced from human-written, post-edited, and synthetic datasets to evaluate models, ensuring coverage of a wide range of domains and linguistic structures (Table 1). In this work, we focus on the monolingual setting of LCB, where the prompt and expected response are in the same language. This setting is particularly relevant for mechanistic interpretability research, as it isolates language confusion phenomena from the additional complexities of explicit cross-lingual transfer.
To quantify language confusion, we adopt two key metrics from LCB: line-level pass rate (LPR) and line-level language accuracy (Acc). LPR measures the percentage of model responses in which every line is in the correct language. Acc reflects the proportion of individual lines across all responses that are correctly generated in the target language. Both metrics rely on automatic language identification using the fastText classifier (Joulin et al., 2016, 2017), which efficiently detects the language of each line in the generated output.
We conducted preliminary benchmarking experiments on LCB with three instruction-tuned LLMs: Llama3-8B (English-centric, no multilingual instruction tuning), Llama3-8B-multilingual (multilingual instruction-tuned) (Devine, 2024), and Llama3.1-8B ( multilingual-optimized). As shown in Figure 1c, Llama3-8B exhibits substantial language confusion, with frequent line-level switches to unintended languages (mostly English). In contrast, both Llama3-8B-multilingual and Llama3.1-8B achieve near-perfect LPR and line-level accuracy, demonstrating the effectiveness of multilingual instruction tuning and targeted optimization for multilingual dialogue.
Given these findings, our work centers on understanding and mitigating the language confusion observed in English-centric Llama3-8B. By leveraging mechanistic interpretability methods, we aim to uncover the internal causes of confusion and develop interventions that can bring its performance closer to that of explicitly multilingual-tuned models. In the following subsection, we delve deeper into the significance of confusion points as critical junctures in the generation process.
3.2 Significance of Confusion Points
A confusion point (CP) is the position in a model’s output where the first token of an unintended language abruptly appears, marking the onset of language confusion (Marchisio et al., 2024). This concept is inspired by psycho- and neurolinguistic research on code-switching, where the precise location of a language switch—known as a switch point—is central to understanding bilingual language production and processing (Blanco-Elorrieta and Pylkkänen, 2017; Suurmeijer et al., 2020). To empirically assess the role of CPs in LLM language confusion, we conduct a replacement experiment on Llama3-8B. For each instance of language confusion, we identify the CP using the fastText language detector. We then replace the token at the CP with the corresponding token generated by Llama3-8B-multilingual, which achieves near-perfect language accuracy, under the same prompt. This approach is motivated by the psycholinguistic observation that, in human code-switching, the choice at the switch point strongly influences the subsequent language trajectory (Moreno et al., 2002; Lai and O’Brien, 2020).
Figure 2: Average token counts and probabilities for English and target language tokens among the top-10 predictions at each layer, shown for both correct and confusion samples across four languages from Aya.
Our results, summarized in Table 2, show a substantial reduction in language confusion after CP replacement, even though our method does not represent an oracle upper bound. These findings highlight the centrality of confusion points in the emergence of language confusion and motivate our subsequent mechanistic analysis and targeted interventions.
4 Mechanistic Analysis of Language Confusion Points
4.1 Analyzing Layer-wise Language Transition
A central question in understanding language confusion is where and how the model’s internal representations fail to transition from a shared conceptual space to the intended target language. Motivated by recent findings that English-centric LLMs process information in a latent, often English-biased, conceptual space before converting it to the target language in the final layers (Wendler et al., 2024; Wang et al., 2025a), we conduct a detailed layer-wise analysis of this transition using TunedLens (Belrose et al., 2023).
We employ TunedLens, the more reliable variant of LogitLens (Nostalgebraist, 2020), to unembed the hidden states of Llama3-8B at each layer into the vocabulary space. With this, we inspect every layer of the model and extract the top 10 predicted tokens with the largest logits at the position immediately preceding the confusion point (CP) (for confusion cases) or the output token (for correct cases). For each layer, we compute the average number and summed probabilities of English and target language tokens among the top-10 predictions, using fastText for language identification. Our analysis focuses on four typologically diverse languages (Arabic, Portuguese, Turkish, Chinese) from the LCB benchmark. We separate samples into two groups: (1) Correct—where the model generates the intended language throughout, and (2) Confusion—where the model switches to an unintended language at a CP. For confusion samples, we analyze the model’s state up to the token before the CP.
Figure 2 presents the evolution of language token counts and probabilities across layers for both groups. In early and middle layers, English tokens dominate the top-10 predictions for all languages, reflecting the English-centric latent conceptual space of Llama3-8B. This is consistent with prior work showing that LLMs encode information in a shared, language-agnostic space in intermediate layers. In the final layers, a sharp transition emerges. For correct samples, the number and probability of target language tokens rise steeply, overtaking English tokens in the last few layers—indicating a successful transition to the target language surface form. In contrast, for confusion samples, this transition fails: English tokens remain dominant or even increase, while target language tokens lag behind. This failure to shift from the latent conceptual space to the target language at the critical moment leads to CPs and erroneous output.
Our layer-wise analysis with TunedLens reveals that the transition to the target language occurs in the final layers, and that failures in this process are tightly linked to language confusion. These findings provide direct evidence that language confusion in Llama3-8B is primarily caused by transition failures in the last few layers, motivating our subsequent neuron-level investigation to pinpoint and intervene on the specific components responsible for these failures.
4.2 Localizing Critical Neurons at Confusion Points
A key step toward understanding and mitigating language confusion is to identify which neurons are most responsible for the emergence of confusion points. Building on recent advances in neuron-level attribution (Geva et al., 2022; Yu and Ananiadou, 2024b), we adopt a static, efficient method to locate and analyze the most influential feed-forward network (FFN) neurons in Llama3-8B.
Methodology
In the inference pass in decoder-only LLMs, for a given input sequence, each layer output $h_{i}^{l}$ (layer $l$, token position $i$) is a sum of the previous layer’s output $h_{i}^{l-1}$, the attention output $A_{i}^{l}$, and the FFN output $F_{i}^{l}$:
$h_{i}^{l}=h_{i}^{l-1}+A_{i}^{l}+F_{i}^{l}$ (1)
The FFN output $F_{i}^{l}$ is calculated by a non-linear $\sigma$ on two MLPs $W^{l}_{fc1}\in\mathbb{R}^{N\times d}$ and $W^{l}_{fc2}\in\mathbb{R}^{d\times N}$:
$F_{i}^{l}=W^{l}_{fc2}\sigma(W^{l}_{fc1}(h_{i}^{l-1}+A_{i}^{l}))$ (2)
Following  Geva et al. (2021), the FFN layer output $F_{i}^{l}$ can be represented as a weighted sum over neuron subvalues:
$F_{i}^{l}=\sum_{k=1}^{N}m_{i,k}^{l}\cdot fc2^{l}_{k}$ (3)
$m_{i,k}^{l}=\sigma(fc1^{l}_{k}\cdot(h_{i}^{l-1}+A_{i}^{l}))$ (4)
where $fc2^{l}_{k}$ is the $k$-th column of $W^{l}_{fc2}$, and $m_{i,k}^{l}$ is derived from the inner product between the residual output $(h_{i}^{l-1}+A_{i}^{l})$ and $fc1^{l}_{k}$, the $k$-th row of $W^{l}_{fc1}$.
Geva et al. (2022) and Dar et al. (2023) project FFN neuron subvalues with unembedding matrices to compute the token probability distribution. To quantify the importance of each neuron for generating a specific token (e.g., at a confusion point), we adopt the log probability increase method of Yu and Ananiadou (2024b). For a neuron in the $l$-th FFN layer $v^{l}$, its importance score is defined as the increase in log probability of the target token when $v^{l}$ is added to the residual stream $A^{l}+h^{l-1}$, compared to the baseline without $v^{l}$:
$\begin{split}Imp(v^{l})=\log(p(w|v^{l}+A^{l}+h^{l-1})-\\ \log(p(w|A^{l}+h^{l-1})\end{split}$ (5)
This approach efficiently identifies neurons whose activations most strongly influence the model’s prediction at a given position.
Experimental Observations
Figure 3: Distribution of Important Neurons Associated with Confusion Points in Llama3-8B. (a) Distribution of the top 300 most important FFN neurons across layers for an individual Chinese prompt “{CJK*}UTF8gbsn请解释拆东墙补西墙的意思。(Please explain ‘{CJK*}UTF8gbsn拆东墙补西墙.’)” from Aya. (b) Aggregated distribution of important neuron scores across all Chinese test samples in Aya.
We apply this method to Llama3-8B on confusion samples from the LCB benchmark, focusing on the token position immediately preceding each confusion point. For each sample and language, we compute the importance scores for all 14,336 FFN neurons in each layer of Llama3-8B, rank them, and select the top 300 most important neurons per sample. We then analyze the distribution of these critical neurons across layers, both for individual samples and aggregated over all samples in a language. Our analysis reveals a striking concentration of important neurons in the final layers, as visualized in Figure 3. This pattern holds both at the single-sample level and when aggregating across samples, indicating that the emergence of confusion points is primarily driven by late-layer FFN activity. We further rank neurons by their frequency of appearance in the top 300 sets across samples, finding that a subset of neurons consistently recurs as highly influential for confusion points.
ar pt tr zh es fr hi ru ja ko de id it vi Avg.
original 33.44 74.26 37.55 24.04 77.15 63.16 16.47 28.20 17.44 23.50 23.00 27.33 39.83 14.79 35.73
freq 31.75 75.10 36.51 22.09 76.29 66.98 18.66 27.70 19.29 23.08 22.25 27.83 39.45 13.58 35.75
score 76.97 93.41 67.61 80.63 91.22 74.77 60.00 50.32 53.50 33.25 40.27 53.58 96.00 67.56 67.08
comparative 85.45 97.12 57.27 89.39 92.20 83.17 82.74 89.43 49.95 40.33 80.82 78.94 95.25 66.50 77.75
Table 3: Confusion mitigation performance of different selection strategies. Line-level accuracy is reported.
To understand the effect of multilingual alignment, we repeat the analysis on Llama3-8B-multilingual using the same set of prompts. After multilingual instruction tuning, language confusion is nearly eliminated. Comparing neuron importance scores between the two models (Figure 4), we observe that most neurons critical for confusion in the Llama3-8B become much less important in its multilingual counterpart, suggesting that multilingual alignment suppresses the activity of confusion-inducing neurons. However, a small number of neurons remain important or even increase in importance, likely reflecting their role in encoding general semantic information rather than language-specific transitions.
Figure 4: Neuron rank comparison between original Llama3 and multilingual Llama3. Results of Chinese test samples in Aya.
These findings reinforce the conclusion from our layer-wise analysis: language confusion is tightly linked to the activity of specific FFN neurons in the final layers. The suppression of these neurons through multilingual alignment provides a mechanistic explanation for the effectiveness of such tuning. Moreover, the identification of a small set of persistent, semantically important neurons suggests that targeted neuron-level interventions could mitigate confusion without harming overall model performance. These insights directly inform our subsequent strategies for neuron-based mitigation of language confusion.
5 Mitigating Language Confusion via Neuron Editing
A central challenge in mitigating language confusion via neuron editing is to identify a set of neurons whose intervention effectively reduces confusion without degrading the model’s general competence or fluency. Insights from our previous mechanistic analysis indicate that language confusion is primarily driven by a subset of late-layer FFN neurons. However, indiscriminate deactivation of important neurons risks harming the model’s overall performance. Thus, a principled neuron selection strategy is essential.
token_num token_prob fluency acc_ood xnli senti
Original 1.96 24.5 25.8 39.9 46.4 98.4
Edited 3.43 36.8 21.8 74.25 44.9 98.2
Diff 1.47 12.3 -4.0 34.4 -1.5 -0.2
Table 4: Results of generalization and robustness of neuron editing. Average performance across languages is reported. Detailed results in Appendix B.
5.1 Neuron Selection and Intervention
We compare three neuron selection strategies: (1) Frequency-Based Selection: Selects the neurons most frequently identified as important across all confusion samples for a given language. (2) Aggregate Importance Selection: Ranks neurons by the sum of their importance scores across all confusion samples, selecting those with the highest cumulative influence. While this method captures the overall impact, it may still include neurons essential for general language competence. (3) Comparative Importance Selection: Inspired by Yu and Ananiadou (2024a), this strategy identifies neurons whose importance scores for confusion points decrease most substantially after multilingual alignment. Specifically, for each neuron, we compute the difference in importance score between original Llama3-8B and Llama3-8B-multilingual on the same input. Neurons with the largest drop are prioritized for intervention, as they are likely to be specifically implicated in language confusion rather than general semantic processing.
For each strategy, we select the top 100 neurons and intervene by setting their activations to zero during generation. We evaluate the impact of each method on the LCB benchmark. Our results (Table 4) demonstrate that Comparative Importance Selection achieves the most effective reduction in language confusion, substantially outperforming both frequency-based and aggregate importance methods. Frequency-based selection yields minimal benefit, while aggregate importance provides moderate improvement but still lags behind our proposed approach. Notably, the comparative strategy selectively targets neurons implicated in confusion, minimizing collateral impact on general model competence.
5.2 Generalization and Robustness of Neuron Editing
To further validate the effectiveness and safety of our Comparative Importance Selection strategy, we conduct a comprehensive evaluation across multiple metrics and experimental setups. Our goal is to ensure that neuron editing not only mitigates language confusion but also preserves the model’s general competence, fluency, and robustness across domains (Table 4).
Language Confusion Mitigation
We first assess the impact of neuron editing on language confusion using the LCB benchmark. In addition to standard metrics (line-level pass rate and line-level accuracy), we analyze the internal output distributions by reporting (1) the number of target language tokens among the top-10 candidates in the final output token logit, and (2) the total probability mass assigned to target language tokens in the top-10. These metrics provide a deeper view of how neuron editing shifts the model’s internal preference toward the intended language, beyond surface-level accuracy.
Robustness on General Tasks
To evaluate whether neuron editing affects the model’s general capabilities, we test the edited model on widely used multilingual benchmarks, including XNLI and multilingual sentiment analysis. We also assess output fluency by measuring the perplexity of generated responses using the multilingual model facebook/xglm-564M (Lin et al., 2022). Across all these metrics, the edited model maintains performance comparable to the original Llama3-8B, indicating that our intervention does not degrade general language understanding or generation quality.
Out-of-Domain Generalization
We further examine the generalization of neuron editing by applying neurons selected from one data source (e.g., Aya) to out-of-domain test sets (e.g., Okapi) for the same language. The edited model continues to demonstrate strong mitigation of language confusion, suggesting that the identified neurons capture robust, domain-independent mechanisms underlying confusion points.
5.3 Comparison with Multilingual Alignment
To contextualize the effectiveness of neuron editing, we compare the performance of the edited Llama3-8B model with that of the multilingual-tuned Llama3-8B. Quantitative results show that neuron editing achieves language confusion mitigation on par with the multilingual-aligned model for most languages. Both approaches yield high line-level pass rates and language accuracy, demonstrating that targeted neuron intervention can match the benefits of extensive multilingual instruction tuning for confusion reduction.
Beyond aggregate metrics, qualitative analysis reveals further advantages of the neuron editing approach. Through case studies, we observe that the edited Llama3-8B not only generates fluent and accurate responses in the intended target language, but also avoids certain artifacts introduced by multilingual alignment. For example, in several instances (Figure 5), the multilingual Llama3-8B produces outputs containing HTML tags or formatting patterns reminiscent of its instruction tuning data, reflecting the influence of imperfect or noisy multilingual datasets. In contrast, the neuron-edited model consistently produces clean, well-structured, and contextually appropriate responses, free from such extraneous formatting.
Figure 5: Case study of neuron editing.
These findings highlight a key strength of mechanistic neuron editing: it directly addresses the internal causes of language confusion without introducing side effects from large-scale data-driven alignment. By preserving the original model’s semantic competence and output quality, neuron editing offers a more targeted and interpretable solution. This suggests that, beyond traditional multilingual instruction tuning, mechanistic interpretability-driven interventions can provide a promising path toward high-quality, robust multilingual language models.
6 Conclusions
This work provides the first mechanistic interpretability account of language confusion in English-centric LLMs. By tracing confusion points to failures in late-layer transitions and localizing the critical neurons responsible, we demonstrate that targeted neuron editing can robustly mitigate language confusion without sacrificing general competence or fluency. Our approach achieves results on par with multilingual-tuned models for most languages, while preserving cleaner output quality. These findings highlight the promise of neuron-level interventions for more reliable and interpretable multilingual language modeling.
Limitations
While this work provides the first mechanistic interpretability account of language confusion in English-centric LLMs, several limitations remain. Our analysis primarily focuses on the monolingual setting; cross-lingual contexts, which may involve distinct mechanisms and challenges, are left for future research. Additionally, neuron editing interventions are evaluated on selected benchmark tasks and may require further validation across broader domains and model architectures. Lastly, while our approach identifies and mitigates language confusion, fully understanding how these mechanisms interact with other multilingual phenomena warrants further investigation.
Ethic Statement
This research was conducted in accordance with the ACM Code of Ethics. The datasets that we use are publicly available. We have not intended or do not intend to share any Personally Identifiable Data with this paper. Regarding the usage of AI tools, we only use AI models for language refining.
Appendix A Further Discussion on Code-Switching and Language Confusion
Code-switching as a Linguistic Phenomenon
Code-switching, the practice of alternating between languages within a single conversation or utterance, is a well-studied natural phenomenon in bilingualism and psycholinguistics (Gardner-Chloros, 2009). Code-switching is typically intentional, often reflecting speakers’ identities, social relationships, and contextual adaptation (Treffers-Daller, 2009; Yim and Clément, 2021). In NLP, code-switching has been explored through evaluating model performance on code-switched data for tasks such as sentiment analysis, machine translation, summarization, and language identification (Khanuja et al., 2020; Doğruöz et al., 2021; Winata et al., 2023). Code-switching is a natural, contextually appropriate strategy in human communication, whereas language confusion, on which our work focuses, is an unintended and erroneous switch to an incorrect language in LLMs (Marchisio et al., 2024). Though related to code-switching, language confusion is an unnatural phenomenon that arises from model failures rather than communicative intent.
Language Confusion and Confusion Points in LLMs
Language confusion has been observed in various multilingual NLP settings, such as “source language hallucinations” in zero-shot cross-lingual transfer (Li and Murray, 2023; Pfeiffer et al., 2023; Chirkova and Nikoulina, 2024) and “off-target translation” in machine translation (Sennrich et al., 2024). In LLMs, this manifests as abrupt, unexpected switches to the wrong language during generation, even under explicit instructions. This issue is particularly prevalent in English-centric models lacking robust multilingual alignment (Zhong et al., 2024). A key concept in recent work is the confusion point—the specific position in generation where the model transitions to an unintended language. Inspired by the importance of code-switching points in human bilingualism, confusion points are central to understanding and diagnosing language confusion in LLMs (Guzzardo Tamargo et al., 2016). Unlike natural code-switching, these points reflect internal model failures. Recent benchmarks (Marchisio et al., 2024) systematically characterize confusion points at response, line, and word levels, revealing their widespread impact and motivating deeper mechanistic investigation, as pursued in this work.
Appendix B Full Experimental Results
Table 5 presents the full benchmarking results. Table 6 shows the full results of the CP replacement experiment. Tables 7 and 8 present the full results of robustness and generalization experiments.
Appendix C Detailed Experimental Setup
C.1 Models
We primarily use three variants of the Llama3 family for our experiments:
  • Llama3-8B: The baseline English-centric model without multilingual instruction tuning.
  • Llama3-8B-multilingual: The multilingual instruction-tuned version, as described in (Devine, 2024).
  • Llama3.1-8B: An improved model optimized for multilingual dialogue.
All models are used in their publicly released forms unless otherwise stated. For neuron editing experiments, we intervene on Llama3-8B using the strategies described in Section 5.
C.2 Datasets and Tasks
Language Confusion Benchmarking and Replacement Experiments
We use the Language Confusion Benchmark (LCB) (Marchisio et al., 2024) for all language confusion detection and mitigation experiments. LCB covers 15 typologically diverse languages and comprises several monolingual and cross-lingual datasets:
  • Monolingual sources: Aya (human-generated), Dolly (post-edited), Native (human-generated), and Okapi (synthetic + machine translated).
  • Languages: Arabic, English, Portuguese, Turkish, Chinese, Spanish, French, Hindi, Russian, Japanese, Korean, German, Indonesian, Italian, Vietnamese.
All main benchmarking and confusion point replacement experiments are run on the monolingual portions of LCB, using 100 prompts per language per dataset as described in Table 1.
Robustness and Generalization Experiments
To assess the robustness and generalization of neuron editing, we evaluate on:
  • XNLI (Conneau et al., 2018): Cross-lingual natural language inference in 15 languages.
  • Multilingual Sentiment Analysis: Standard multilingual sentiment datasets (including German, Spanish, French, Japanese, and Chinese). It is a binary classification task derived from the multilingual Amazon review dataset.
  • Out-of-domain LCB evaluation: For each language, neurons are selected from one LCB source (e.g., Aya), then tested on a different source (e.g., Okapi) to assess generalization.
C.3 Metrics
Language Confusion Metrics
We adopt two primary metrics from LCB:
  • Line-level Pass Rate (LPR): Percentage of responses where every line is in the correct language.
  • Line-level Accuracy: Proportion of lines generated in the correct language.
Language identification for these metrics is performed using the fastText classifier (Joulin et al., 2016).
Internal Model Metrics
We further report:
  • Target Language Token Count: Number of target language tokens among the top-10 output logits in the final layer.
  • Target Language Token Probability: Total probability mass assigned to target language tokens in the top-10 output logits.
Generalization and Fluency Metrics
  • XNLI and Sentiment Accuracy: Standard classification accuracy on XNLI and multilingual sentiment analysis tasks.
  • Fluency (Perplexity): Perplexity of generated outputs, measured using the multilingual facebook/xglm-564M model (Lin et al., 2022).
C.4 Implementation Details
All experiments are run on NVIDIA A100 GPUs. Prompt formatting and decoding settings follow the LCB benchmark defaults. Neuron interventions are implemented at inference time via custom hooks in PyTorch, zeroing out selected neuron activations layer-wise as described in Section 5.1. For TunedLens analysis, we use the public implementation from Belrose et al. (2023).
All code, evaluation scripts, and neuron selection details will be released upon publication to facilitate reproducibility.
Table 5: Full benchmarking results on LCB.
Table 6: Full results of CP replacement experiments
num_ori prob_ori num_edit prob_edit num_diff prob_diff fluency_ori fluency_cna diff
ar 2.83 25.8 5.37 30.3 2.55 4.5 30.1 24.7 -5.4
pt 2.86 49.5 3.41 56.0 0.56 6.5 25.7 23.3 -2.3
tr 2.05 29.5 2.42 23.5 0.37 -6.0 21.2 18.8 -2.5
zh 1.33 8.6 5.10 37.3 3.78 28.7 33.1 26.0 -7.0
es 1.67 26.5 3.28 50.3 1.61 23.8 25.4 23.2 -2.2
fr 2.48 43.0 2.91 49.2 0.43 6.2 21.2 21.1 -0.1
hi 1.25 12.0 1.64 13.7 0.39 1.8 28.5 22.9 -5.6
ru 1.09 18.0 3.21 31.0 2.12 13.0 23.7 19.5 -4.2
de 2.73 23.7 4.45 37.1 1.72 13.4 23.8 18.5 -5.3
it 1.33 8.4 2.50 39.3 1.17 31.0 25.7 20.2 -5.5
avg 1.96 24.5 3.43 36.8 1.47 12.3 25.8 21.8 -4.0
Table 7: Full results of robustness experiments. Perplexity is calculated to measure fluency.
xnli
language acc_ori acc_edit
ar 0.42 0.37
de 0.54 0.54
es 0.46 0.5
fr 0.49 0.5
hi 0.47 0.48
ru 0.37 0.3
tr 0.46 0.52
vi 0.46 0.37
zh 0.51 0.46
avg 0.464 0.449
sentiment analysis
language acc_ori acc_edit
de 0.98 0.98
es 0.98 0.98
fr 0.98 0.97
ja 0.99 0.99
zh 0.99 0.99
avg 0.984 0.982
Table 8: Full results of generalization experiments.